OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (8,-4) ; Slope: -5/6

OpenStudy (anonymous):

Point slope form is y = mx + b m = slope This can let us have y = -5x/6 + b In order to find b, you have to substitute y and x into the equation -4 = -5(8)/6 + b -4 = -40/6 + b -4 = -20/3 + b -12/3 + 20/3 = b 8/3 = b Does that make sense?

OpenStudy (anonymous):

y = -5x/6 + 8/3

OpenStudy (anonymous):

Hmm, there are only answer choices A. y+4=-5/6(x+8)

OpenStudy (anonymous):

B. y-4=-5/6(x+8)

OpenStudy (anonymous):

C. y+4=-5/6(x-8)

OpenStudy (anonymous):

D. y-4=5/6(x-8)

OpenStudy (anonymous):

The answer is C Point-slope form is: y-y1=m(x-x1). So y+4=-5/6(x-8)